OpenStudy (anonymous):

find the area of the region between the graph of y=(e^x) sinx and x-axis on the interval [0,Pi].

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi}e ^{x}\sin x dx\] now do math :)

OpenStudy (anonymous):

oh yes i know it lol then can u solve it ?

OpenStudy (anonymous):

\[(e ^{\pi}+1)/2\] (i think) See method in http://en.wikipedia.org/wiki/Integration_by_parts#Liate_rule

OpenStudy (anonymous):

ok thx i ll check and try to solve